if sina+cosa=underoot2cosa then cosa-sina is equal to(a=theta)

Grade:12

1 Answers

Ramesh V
14 years ago

sin A = (21/2 - 1) cos A

tan A = (21/2 - 1)

cos A - sin A = (2 - 21/2) / (4 - 2.21/2)1/2

                  = (21/2 - 1)1/2

--

regards

Ramesh